OpenStudy (anonymous):

-(1+7x)-6(-7-x)=36 and show work, please?

OpenStudy (anonymous):

-(1+7x)+6(7+x)=36 -1-7x+42+6x=36 -x+41=36 x=5

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome